OpenStudy (anonymous):

Given matrix D, which matrix represents -4D? D= -4 8 -12 4

OpenStudy (anonymous):

multiply each element within the matrix by -4 : \[\left[\begin{matrix}-4*-4 & 8*-4 \\ -12*-4 & 4*-4\end{matrix}\right]=\left[\begin{matrix}(?) & (?) \\ (?) & (?)\end{matrix}\right]\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome